OpenStudy (shuckle):

Write each quotient as a complex number. -2i / 1+i Is -2/-1 and +2i/-1 the correct answer?

zepdrix (zepdrix):

\[\large\rm \frac{-2i}{1+i}\left(\frac{1-i}{1-i}\right)=\frac{-2i-2}{2}\]Hmm, do you understand this step that I applied?

OpenStudy (shuckle):

Yes, I understand that you have to multiply the numerator and denominator by the complex conjugate of the denominator. But from doing that I got the numbers\[\frac{ -2+2i^{2}}{ 1-1i+1i-1^{2}}\]

OpenStudy (shuckle):

Wow that thing is hard to use. -_-

zepdrix (zepdrix):

Woops your denominator is \(\large\rm 1-1i+1i-i^2\)

OpenStudy (shuckle):

I don't understand?

OpenStudy (shuckle):

Oh I see, the one is an i. But I still don't know what the final solution should be?

zepdrix (zepdrix):

So the middle terms cancel out in the denom, ya? :)\[\large\rm \frac{ -2+2i^{2}}{ 1-1i+1i-i^{2}}=\frac{ -2+2i^{2}}{ 1-i^{2}}\]

zepdrix (zepdrix):

i^2 is -1, so the negatives give us a +1 in that spot

OpenStudy (shuckle):

Yes, and then multiplied to the denominator of +1 give you a negative one on the bottom?

zepdrix (zepdrix):

\[\large\rm =\frac{ -2+2(-1)}{ 1-(-1)}\]

zepdrix (zepdrix):

No you're not multiplying, you're adding down there.

OpenStudy (shuckle):

Oh.

zepdrix (zepdrix):

Oh woops that's a 2i for the first term in the numerator :O

zepdrix (zepdrix):

\[\large\rm =\frac{ -2i+2(-1)}{ 1-(-1)}\]

OpenStudy (shuckle):

So after that it should be -2i -2 over 1?

zepdrix (zepdrix):

No, your denominator simplifies to this:\[\large\rm =\frac{ -2i+2(-1)}{ 1+1}\]

OpenStudy (shuckle):

because (-1)(-) makes a positive one, right?

zepdrix (zepdrix):

ya :)

OpenStudy (shuckle):

So then, it'd be -2+2i over 1+1 which will equal -2 +2i over 2, and then I should divide?

OpenStudy (shuckle):

So the answer is -1-i?

zepdrix (zepdrix):

yayyy good job \c:/

OpenStudy (shuckle):

Thank you so much. :3